OpenStudy (anonymous):

Teesha is in French club. There are 19 students in the club. The French teacher will pick two students at random to guide visiting students from France. What is the probability that Teesha will not be picked up as a guide? a. 2/19 b. 17/19 c. 1/18 d. 18/19

OpenStudy (anonymous):

@RainbowsxD ??

OpenStudy (anonymous):

@ayyookyndall you kno?

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

are you sure these are the possible answers?

OpenStudy (anonymous):

thats what was on my review? @misty1212

OpenStudy (misty1212):

oh ok i see it

OpenStudy (misty1212):

the probability she is not picked on the first try is \[\frac{18}{19}\] then not picked on the second try given that she is not picked on the first try is \[\frac{17}{18}\] multiply these together and get \[\frac{18}{19}\times \frac{17}{18}=\frac{17}{19}\]

OpenStudy (anonymous):

oh ok that makes more sense! thanks a lot!(: @misty1212